# Principled Federated Random Forests for Heterogeneous Data

### v1

@@ 4.3 FedForest Training Protocol
22:Finalize: Server computes leaf values from final node summaries.
Communication per node. We count communication in transmitted scalars for one node $\nu$ and one client $k$, with $d$ active features, $B$ quantile levels (sketch size), and sufficient-statistic size $S$ (e.g., $S=3$ for regression and $S=C$ for $C$-class classification). All costs below are independent of the node sample sizes $(n_{\nu,k},n_{\nu})$.
FedForest (exact gain). Communication has two steps. (i) Sketching: the client sends $d$ univariate quantile sketches of size $B$, for a cost $dB$. (ii) Evaluation: for each tested threshold in $\mathcal{T}$, the client returns the left-child sufficient statistics, costing $S$ scalars per candidate. So the total cost of FedForest per node is $O(dBS)$.
Top-$L$ shortlisting. If the server evaluates only $|\mathcal{L}_{\nu}|\ll d$ shortlisted features, the evaluation term becomes $|\mathcal{L}_{\nu}|BS$.
AvgImp (Top-$L$ + averaged local gains). If clients compute their impurity gains locally at the candidate thresholds of the shortlisted features, they send one scalar per threshold instead of $S$. Combining Top-$L$ with AvgImp therefore reduces the evaluation communication to $|\mathcal{L}_{\nu}|B$, at the price of shifting computation from the server to the clients for the local gain evaluation.
5 Experiments
5.1 Synthetic Data
Experimental setting. We simulate an horizontal multicenter setting with $K$ clients, where the $k$-th client holds $n_{k}$ i.i.d. samples $(H_{i}\!=\!k,X_{i}^{(k)},Y_{i}^{(k)})$. Covariates are drawn from client-specific Gaussians $X_{i}^{(k)}\sim\mathcal{N}(\mu_{k},\alpha_{k}I_{d})$, with $\mu_{k}=[(-1)^{k}\gamma,\dots,(-1)^{k}\gamma]$, $\gamma\geq 0$ and $\alpha_{k}>0$. The separation between client distribution can be tuned via $\gamma$ and $\alpha_{k}$. Choosing a small values for $\gamma$ and large values for $\alpha_{k}$ corresponds to client distributions with large overlaps. Observed outcomes are generated as $Y_{i}^{(k)}=f(X_{i}^{(k)})+\delta_{k}+\varepsilon_{i}$, where $\varepsilon_{i}\sim\mathcal{N}(0,\sigma^{2})$ is an independent noise, $\delta_{k}=(-1)^{k}\delta$ is a client-specific shift controlling outcome heterogeneity, where the function $f$ is a decision-tree function (see Appendix D). We compare our proposed methods to several federated baselines and a centralized RF, considered as the reference as it is trained on pooled data (via scikit-learn, see Pedregosa et al., 2011). Federated baselines include:
  1. 1.
    Federated Histogram (Kalloori and Klingler, 2022), adapted to regression via node-wise sufficient statistics for squared error (counts, sums, sums of squares);
  2. 2.
    One-shot local ensembling (Hauschild et al., 2022), where clients train local forests and the server aggregates predictions;
  3. 3.
    Local learning, where each client trains on $\mathcal{D}_{k}$ only.
We compare these competitors to our proposed methods:
  1. 1.
    FedForest-Quantiles, which uses quantile sketches to propose candidate thresholds and exact split evaluation via additive sufficient statistics;
  2. 2.
    FedForest-AvgImp/Top-$L$, our cheaper variant that shortlists features using Top-$L$ and evaluates candidates using averaged local impurity gains.
All RFs use $T=50$ trees, with a maximum depth of $8$, a minimum leaf size of $5$, feature subsampling $\sqrt{d}$, and client-stratified bootstrapping. For candidate generation, FedForest methods and Federated Histogram use $B=32$ candidates per feature, and the FedForest AvgImp method screens $L=3$ features.
In our methods, $H$ is always included as a split candidate (i.e., not subject to feature subsampling) so that the model can capture client effects when they reduce impurity. Baselines that do not implement client splits are evaluated on $X$ only. Whereas not natively supported for regression RFs, using Fisher’s encoding of the client indicator improves the predictive performance of centralized RFs that allow splits on $H$. We use such encoding at the root node for centralized RFs, while our method recomputes this encoding at every node based on the shared node summaries.
We consider $K=10$ clients holding $d=20$ features, and $n_{k}=200$ samples per client, with the three settings:
  1. 1.
    (Homogeneous) $\forall k\in[K],\gamma_{k}=\delta=0$ and $\alpha_{k}=1$;
  2. 2.
    (Covariate shift) $\forall k\in[K]$, $\gamma_{k}=3$ and $\alpha_{k}=0.5$;
  3. 3.
    (Outcome shift) $\delta=1.5$, $\forall k\in[K],\alpha_{k}=1$ and $\gamma_{k}=0$ (homogeneous covariates).
Figure 2: Methods comparison on homogeneous clients
Figure 3: Methods comparison under covariate shift
Figure 4: Methods comparison under outcome shift
Results. Figure 2 reports, in the homogeneous setting, the test mean squared error (MSE) on a $30\%$ held-out set drawn from the client mixture, averaged over $150$ synthetic data generations. Here, our proposed FedForest Quantiles (regardless of including $H$ as a splitting variable) and Federated Histogram (which does not use $H$ for splitting) are comparable to the centralized RF. FedForest without including $H$ exhibits a lower variance than Federated Histogram. Local methods (either local learning or local ensembling), which are built independently on each client, perform poorly, due to the smaller effective sample size per client. Indeed, while no bias is expected from local learning in this setting, each tree is trained on less data compared to other federated or centralized methods. Our approximate variant FedForest AvgImp (averaging impurity gains) performs slightly less well, and worsens when $H$ is included; with larger candidate grids $B$ or more screened features $L$, it reaches centralized RF performance.
In the covariate shift scenario (Figure 3), the chosen $(\gamma_{k},\alpha_{k})$ yield a shift close to disjoint client supports, and we design the outcome to require a split near 0 between these supports, making this an extreme regime that clearly exposes flawed procedures. As in the homogeneous case, FedForest Quantiles (with or without $H$) and Federated Histograms are competitive with centralized RF. FedForest AvgImp performs poorly, as expected outside its intended (approximately i.i.d.) regime: with disjoint supports, averaging clientwise impurity gains can miss splits between supports because each client’s impurity reduction is then zero (all its samples fall on one side). Allowing AvgImp to split on $H$ improves performance since $H$ acts as a proxy for identifying the support of the client distribution. On the contrary, methods that select the best split based on an impurity reduction computed with a mixture of all client data (as centralized RF or the three federated methods mentioned above) succeed in detecting relevant splits, even when $H$ is not added as a covariate, thus having good predictive performances. Local ensembling is not reported due to its extremely high MSE ($\approx 141.2$), reflecting its inability to handle covariate shift. When client data supports are disjoint, trees are trained independently on each client, and the random forest averages predictions from many irrelevant trees, as a tree trained on one client typically produces near-constant predictions on data from other clients.
Figure 5: Methods comparison on FedHeart-Disease dataset
Finally, Figure 4 (outcome shift) highlights the benefit of client-aware structure: including $H$ substantially improves performance for methods that can split on $H$ (centralized and our FedForest variants), whereas baselines restricted to $X$ cannot capture client effects and incur a larger error. Note that Federated Histograms are not designed to split on $H$ so they fail in this context.
5.2 Real Data
@@
We further validate our approach on the FedHeart Disease dataset of the Flamby federated benchmark (Ogier du Terrail et al., 2022), a binary classification task involving $740$ patients and $13$ clinical covariates distributed across four institutions in the USA (Cleveland, Long Beach), Switzerland, and Hungary. To provide a comprehensive benchmark, we compare our method against two standard federated baselines established in Ogier du Terrail et al. (2022): (i) FedAvg LR, a logistic regression model trained collaboratively with FedAvg (McMahan et al., 2017); and (ii) FedAvg NN, a non-linear two-layer neural network trained via Federated Averaging. Notably, we observe that our methods FedForest Quantiles and FedForest AvgImp have good predictive performances, together with Federated Histograms and Local Learning. Neural networks and Logistic regression fail here. This highlights that Federated Random Forests (FedForest, Federated Histograms, or local learning) outperform these federated parametric models, highlighting the critical importance of flexible, non-linear modeling in this heterogeneous regime.

### v2

@@ 4 Federated Forests
19:Finalize: Server computes leaf values from final node summaries.
Client subsampling (Optional). For a given tree, the server can randomly sample a client subset $\mathcal{K}^{\prime}\subseteq[K]$ (Alg. 1, Line 4) to reduce communication. We recommend this mainly as an acceleration knob under approximately homogeneous clients. Under covariate shift, it changes the mixture distribution and may alter split decisions; under outcome shift, it can omit informative values of $H$.
Communication and computation costs. We summarize and compare the communication and complexities of different methods in Table 1, and derive the per-node scalar costs in Appendix A.
5 Numerical Results
Experimental setting. We consider a centralized scikit-learn Random Forest trained on pooled data as a reference, as well as two federated baselines reimplemented due to the lack of public code: Federated Histogram (Kalloori and Klingler, 2022), adapted to regression via node-wise sufficient statistics for squared loss, and one-shot Local Ensembling (Hauschild et al., 2022), where clients train local forests and the server aggregates predictions. We also include Local Learning, where each client trains a forest on its own dataset $\mathcal{D}_{k}$. We compare these with two variants of our approach: FedForest-Quantiles, which combines quantile sketches with exact split evaluation via additive sufficient statistics, and FedForest-AvgImp/Top-$L$, which uses Top-$L$ feature shortlisting followed by impurity evaluation based on averaged local gains.
All methods use $T=50$ trees, maximum depth $8$, minimum leaf size $5$, feature subsampling ($d/3$ for regression, $\sqrt{d}$ for classification), and client-stratified bootstrapping. FedForest and Federated Histogram use $B=32$ candidates per feature, while AvgImp/Top-$L$ uses $L=3$ screened features. For methods that split on both $X$ and $H$, $H$ is always included as a candidate (i.e., exempt from feature subsampling), whereas baselines without client splits use only $X$. Since standard regression forests do not natively handle unordered categorical splits on $H$, centralized RF uses Fisher encoding at the root node, while FedForest recomputes it at each node from shared summaries.
Synthetic data. We simulate a horizontal multicenter setting with $K=10$ clients. Client $k$ observes $n_{k}=200$ i.i.d. samples $(H_{i}=k,X_{i}^{(k)},Y_{i}^{(k)})$, where $X_{i}^{(k)}\sim\mathcal{N}(\mu_{k},\alpha_{k}I_{d})$ and $\mu_{k}=[(-1)^{k}\gamma,\dots,(-1)^{k}\gamma]$. Here, $\gamma\geq 0$ and $\alpha_{k}>0$ control covariate heterogeneity: smaller $\gamma$ and larger $\alpha_{k}$ induce stronger overlap across clients. Outcomes are generated as $Y_{i}^{(k)}=f(X_{i}^{(k)})+\delta_{k}+\varepsilon_{i}$, with $\varepsilon_{i}\sim\mathcal{N}(0,.5)$ and $\delta_{k}=(-1)^{k}\delta$. Thus, $\delta$ controls outcome heterogeneity, while $f$ is a tree-based regression function (see Appendix F). We set $d=20$ and study three settings: (i) homogeneous ($\gamma=\delta=0$, $\alpha_{k}=1$), (ii) covariate shift ($\gamma=3$, $\alpha_{k}=0.5$, $\delta=0$), and (iii) outcome shift ($\gamma=0$, $\alpha_{k}=1$, $\delta=1.5$).
(a) Homogeneous clients
(b) Covariate shift
(c) Outcome shift
(d) FedHeart-Disease dataset
Figure 2: Performance comparison on simulated data and the real-world FedHeart dataset.
Figure 2(a) reports test MSE on a $30\%$ held-out mixture sample from the client mixture in the homogeneous setting (averaged over 150 draws). FedForest-Quantiles (with or without $H$) and Federated Histogram (which does not split on $H$) perform on par with centralized RF, with FedForest-Quantiles showing lower variance. Local learning and local ensembling perform worse due to limited per-client sample sizes; although unbiased here, they underutilize data relative to federated and centralized methods. FedForest-AvgImp is slightly weaker and degrades when including $H$, but approaches centralized performance with larger $B$ or higher $L$.
In the covariate-shift setting (Figure 2(b)), client supports are nearly disjoint, and the target function requires a split between them. This stresses methods relying on local impurity estimates. FedForest-Quantiles and Federated Histogram remain close to centralized RF since they evaluate splits using pooled statistics. In contrast, FedForest-AvgImp fails in this regime: averaging client-wise gains can miss global splits when each client observes only one side of the partition. Allowing AvgImp to split on $H$ mitigates this by acting as a proxy for client identity. Local ensembling is omitted due to extremely high error ($\approx 141.2$), as models trained per client extrapolate poorly across disjoint supports.
Finally, Figure 2(c) shows the outcome-shift setting. Allowing splits on $H$ significantly improves centralized RF and FedForest variants, which can adapt to client-specific offsets, whereas baselines restricted to $X$ cannot capture this heterogeneity and perform substantially worse. Federated Histogram, not designed to split on $H$, is particularly affected in this regime.
@@ 5 Numerical Results
Real data. We further validate our approach on the FedHeart Disease dataset from the FLamby federated benchmark Ogier du Terrail et al. (2022). This binary classification task comprises $740$ patients and $13$ clinical covariates distributed across four institutions. We compare our approach against two standard federated baselines evaluated by Ogier du Terrail et al. (2022): FedAvg LR, a logistic regression model trained via Federated Averaging (McMahan et al., 2017), and FedAvg NN, a two-layer neural network trained using the same protocol. As shown in Figure 2(d), both FedForest Quantiles and FedForest AvgImp achieve strong predictive performance, on par with the tree-based baselines (Federated Histograms and Local Learning). In contrast, the parametric FedAvg baselines (logistic regression and neural network) perform substantially worse. These results illustrate how flexible tree-based methods are often better suited for heterogeneous tabular datasets.
We further validate our findings on 13 public datasets from OpenML and scikit-learn in Appendix B.